# Equivariant Machine Learning on Graphs with Nonlinear Spectral Filters

### v1

@@ A.3 Stable Invertibility of Synthesis
In the fixed graph setting, given any signal ${\mathbf{X}}$ such that $g_{j}({\bm{\Delta}}){\mathbf{X}}\neq 0$ for every $j\in[K]$, the synthesis operator $\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}$ is stably invertible. Indeed, the right singular vectors of $\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}$ are the standard basis elements $\mathbf{e}_{j}$ in ${\mathbb{R}}^{K}$, the left singular vectors are
@@
$\mathbf{Y}_{j}=\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}}\in{\mathbb{R}}^{N},$
and the singular values are
@@
$\sigma_{j}=\frac{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}},$
for $j\in[K]$. Hence, we have
@@
$\left\lVert\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}\right\rVert_{2}=\max_{j}\sigma_{j}$
@@
$\left\lVert\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{-1}\right\rVert_{2}=\frac{1}{\min_{j}\sigma_{j}}$
Suppose that $a=1$ and $e=0$. In this case, $\mathcal{S}$ is an isometry from the spectral domain to a subspace of the signal space ${\mathbb{R}}^{N}$. Analysis is the adjoint of synthesis.
This analysis can be extended to the index parametrization case.
A.4 Uniform Approximation of Graphs with $J$ Eigenvectors
@@ A.4.1 Cut Norm
The cut norm has the following interpretation, which has precise formulation in terms of the weak regularity lemma [Frieze and Kannan, 1999, Lovász and Szegedy, 2007]. Any pair of (deterministic) graphs are close to each other in cut norm if and only if they can be described as pseudo-random graphs sampled from the same stochastic block model. Hence, the cut norm is a meaningful notion of graph similarity for practical graph machine learning, where graphs are noisy and can represent the same underlying phenomenon even if they have different sizes and topologies. In addition, the distance between non-isomorphic graphons is always positive in cut norm [Lovász, 2012]. In this context, the work in [Levie, 2023] showed that GNNs with normalized sum aggregation cannot separate graphs that have zero distance in the cut norm. This means that the cut norm is sufficiently discriminative for practical machine learning on graphs.
A.4.2 The Constructive Weak Regularity Lemma in Hilbert Spaces
The following lemma, called the constructive weak regularity lemma in Hilbert spaces, was proven in [Finkelshtein et al., 2024]. It is an extension of the classical respective result from [Lovász and Szegedy, 2007].
@@ A.4.2 The Constructive Weak Regularity Lemma in Hilbert Spaces
Lemma A.3.
[[Finkelshtein et al., 2024]] Let $\{\mathcal{K}_{j}\}_{j\in\mathbb{N}}$ be a sequence of nonempty subsets of a real Hilbert space $\mathcal{H}$ and let $\delta\geq 0$. Let $J>0$, let $R\geq 1$ such that $J/R\in\mathbb{N}$, and let $g\in\mathcal{H}$. Let $m$ be randomly uniformly sampled from $[J]$. Then, in probability $1-\frac{1}{R}$ (with respect to the choice of $m$), any vector of the form
@@
Let us use Lemma A.3, with $\mathcal{H}={\mathbb{R}}^{N\times N}$, and $\mathcal{K}_{j}=\mathcal{K}$ the set of symmetric rank one matrices of the form $\mathbf{v}\mathbf{v}^{\top}$ where $\mathbf{v}\in{\mathbb{R}}^{N}$ is a column vector. Denote by $\mathcal{Y}_{m}$ the space of linear combinations of $m$ elements of $\mathcal{K}$, which is the space of symmetric matrices of rank bounded by $m$. For the Hilbert space norm, we take the Frobenius norm. In the setting of the lemma, we take $g=\mathbf{M}$, and $g^{*}\in\mathcal{Y}_{m}$, and $\delta=0$. By the lemma, with probability $1-1/R$, any Frobenius minimizer $\mathbf{C}$, namely, that satisfies $\left\lVert\mathbf{M}-\mathbf{C}\right\rVert_{\rm F}=\min_{\mathbf{C}^{\prime}\in\mathcal{Y}_{m}}\left\lVert\mathbf{M}-\mathbf{C}^{\prime}\right\rVert_{\rm F}$, also satisfies
@@
$\leq$ $\left\lVert\frac{1}{2}\left\langle\bm{\text{1}}_{S\cup T}\bm{\text{1}}_{S\cup T}^{\top},\mathbf{M}-\mathbf{C}\right\rangle\right\rVert_{\rm F}+\left\lVert\frac{1}{2}\left\langle\bm{\text{1}}_{S}\bm{\text{1}}_{S}^{\top},\mathbf{M}-\mathbf{C}\right\rangle\right\rVert_{\rm F}+\left\lVert\frac{1}{2}\left\langle\bm{\text{1}}_{T}\bm{\text{1}}_{T}^{\top},\mathbf{M}-\mathbf{C}\right\rangle\right\rVert_{\rm F}$
@@
where for a set $S\subset[N]$, denote by $\bm{\text{1}}_{S}\in{\mathbb{R}}^{N}$ the column vector with $1$ at coordinates in $S$ and 00 otherwise.
@@
$\left\lVert\mathbf{M}-\mathbf{C}\right\rVert_{\square}\leq\frac{3\alpha}{2}\sqrt{\frac{R}{J}}$
Lastly, note that by the best rank-$m$ approximation theorem (Eckart–Young–Mirsky Theorem [Trefethen and Bau, 2022, Thm. 5.9]), any Frobenius minimizer $\mathbf{C}$ is the projection upon the $m$ leading eigenvectors of $\mathbf{M}$ (or some choice of these eigenvectors in case of multiplicity higher than 1).
@@
Now, if we want to apply Thm. 4.1 to other GSOs ${\bm{\Delta}}$, we need to make some assumptions. Note that when the GSO is a Laplacian, we take as the leading eigenvectors the ones correspoding to the smallest eigenvalues, not the largest ones. To make the theorem applicable, we need to reorder the eigenvectors of ${\bm{\Delta}}$. This can be achieved by applying a decreasing function $h$ to ${\bm{\Delta}}$, such as $h({\bm{\Delta}})={\bm{\Delta}}^{-1}$. The role of $h$ is to amplify the eigenspaces of ${\bm{\Delta}}$ in which most of the energy of signals interest (the ones that often appear in the dataset) is concentrated. Under the assumption that $g({\bm{\Delta}})$ has entries bounded by some not-too-high $\alpha>0$, one can now justify approximating GSOs by low-rank approximations based on the smallest eigenvectors.
A.5 NLSFs on Random Geometric Graphs
@@
We follow a similar analysis to [Paolino et al., 2023], mostly skipping the rigorous Monte-Carlo error rate proofs, as these are standard. Let $\mathcal{S}$ be a metric space with metric ${\rm d}$ and with a Borel probability measure, that we call the uniform measure $dx$. Let $r>0$, and denote by $B_{r}(x)$ the ball of radius $r$ about $x\in\mathcal{S}$. Let $V(x)$ be the volume of $B_{r}(x)$ with respect to the uniform measure. We consider an underlying Laplacian on the space $\mathcal{S}$ defined on signals $f:\mathcal{S}\rightarrow\mathbb{R}$ by
@@
where we assume $C=1$ in the analysis below without loss of generality. Note that $\mathcal{L}$ is compact, so it has a discrete spectrum.
@@
We consider the leading $M$ eigenvectors of $\mathbf{L}$ and $\mathcal{L}$, where $M\ll N$, namely, the eigenvectors corresponding to the largest eigenvalues in their absolute value. We order the eigenvalues in decreasing order from positive to negative. Let $\mathcal{PW}(M)$ be the space spanned by the $M$ leading eigenvectors of $\mathcal{L}$, called the Paley-Wiener space of $\mathcal{L}$. Let $p_{M}$ be the projection upon the Paley-Wiener space of $\mathcal{L}$. Let $R_{N}:L^{2}(\mathcal{S})\rightarrow L^{2}[N]$ be the sampling operator, defined by $(R_{N}f)_{j}=f(x_{j})$.
Let $f:\mathcal{S}\rightarrow{\mathbb{R}}$ be a bounded signal over the metric space and suppose that $f\in\mathcal{PW}(N)$. Denote $\mathbf{f}=(f(x_{n}))_{n=1}^{N}=R_{N}f$. By Monte Carlo theory, we have
@@
So, in case the sampling is uniform. i.e. $w(x)=1$, we have
$L\approx N\mathcal{L},$
pointwise. To make this precise, let us recall Hoeffding’s Inequality.
Theorem A.1 (Hoeffding’s Inequality).
@@
$\mathbb{P}\Big{(}\Big{|}\frac{1}{N}\sum_{n=1}^{N}(Y_{n}-\mathbb{E}[Y_{n}])\Big{|}\geq k\Big{)}\leq 2\exp\Big{(}-\frac{2k^{2}N}{(b-a)^{2}}\Big{)}.$
@@
$\left\lVert R_{N}\mathcal{L}f-\frac{1}{N}{\mathbf{L}}R_{N}f\right\rVert={\mathcal{O}}\Big{(}M\sqrt{\frac{\log(1/p)+\log(N)+\log(2)}{N}}\Big{)}$
@@
We now evoke the transferability theorem – a slight reformulation of Section 3 of Theorem 4 in [Levie et al., 2021].
Theorem A.2.
Let $\mathcal{L}$ be a compact operator on $L^{2}(\mathcal{S})$ and $\mathbf{L}$ an operator on $\mathbb{C}^{N}$. Let $M\in\mathbb{N}$ and let $R_{M}$ be the sampling operator defined above. Let $\mathcal{PW}(M)$ be the space spanned by the $M$ leading eigenvectors of $\mathcal{L}$, and let $f\in\mathcal{PW}(M)$. Let $g:{\mathbb{R}}\rightarrow{\mathbb{R}}$ be Lipschitz with Lipschitz constant $D$. Then, there exists a constant $C$ (that depends on $M$) such that
$\left\lVert R_{N}g(\mathcal{L})f-\frac{1}{N}g({\mathbf{L}})R_{N}f\right\rVert\leq D\sqrt{M}\left\lVert R_{N}\mathcal{L}f-\frac{1}{N}{\mathbf{L}}R_{N}f\right\rVert.$
For every leading eigenvalue $\lambda_{j}$ of $\mathcal{L}$ let $\gamma_{i_{j}}$ be the leading eigenvalue of $\mathcal{L}$ closest to $\lambda$, where, if there are repeated eigenvalues, $i_{j}$ is chosen arbitrarily out of the eigenvalues of $\mathbf{L}$ that best approximate $\lambda_{i}$.Let
$\delta=\min\{\alpha,\beta\}$
@@
$\alpha=\min_{j\in[M]}\min\big{\{}\ \min_{i_{j}\neq i\in[N]}\left\lvert\lambda_{j}-\gamma_{i}\right\rvert\ ,\ \min_{j\neq m\in[M]}\left\lvert\lambda_{j}-\lambda_{m}\right\rvert\ ,\ \min_{j\neq m\in[M]}\left\lvert\gamma_{i_{j}}-\lambda_{m}\right\rvert\ \big{\}}$
@@
$\beta=\min_{i\in[M],n\in[N]}\left\lvert\gamma_{i}-\gamma_{n}\right\rvert$
@@
$g_{j}(\mathcal{L})f=p_{j}^{\mathcal{L}}f,\quad g_{j}(\mathbf{L})R_{N}f=p_{i_{j}}^{\mathbf{L}}\mathbf{f}$
@@
Now, be the transferability theorem,
$\left\lVert R_{N}p^{\mathcal{L}}_{j}f-p^{\mathbf{L}}_{i_{j}}\mathbf{f}\right\rVert\leq D\sqrt{M}\left\lVert R_{N}\mathcal{L}f-\frac{1}{N}{\mathbf{L}}\mathbf{f}\right\rVert={\mathcal{O}}(\sqrt{M/N}).$
By induction over $j$, with base $j=1$, we must have $i_{j}=j$ for every $j\in[M]$.
@@
$\left\lvert\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}^{2}-\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}^{2}\right\rvert={\mathcal{O}}(M^{-1/2})$
@@
and $\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}+\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}$ is bounded from below by the constant $\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}$,
$\left\lvert\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}-\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}\right\rvert={\mathcal{O}}(M^{-1/2})$
@@
Appendix B Additional Details on Experimental Study
We describe the experimental setups and additional details of our experiments in Sec. 5.
B.1 Semi-Supervised Node Classification

### v2

@@ D.3 Stable Invertibility of Synthesis
We present the analysis for diagonal synthesis defined from App. C.1. In the fixed graph setting, given any 1D signal ${\mathbf{X}}$ such that $g_{j}({\bm{\Delta}}){\mathbf{X}}\neq 0$ for every $j\in[K+1]$, we show that the synthesis operator $\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{\text{val, diag}}$ is stably invertible.
Since we consider 1D signal ${\mathbf{X}}$, the diagonal synthesis in filter bank case in Eq. (10) can be written as
$\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{\text{val, diag}}=\mathbf{H}\mathbf{r}^{\prime},$
where
@@
${\mathbb{R}}^{N\times(K+1)}\ni\mathbf{H}=\left[\frac{g_{1}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{1}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}^{a}+e},\ldots,\frac{g_{K+1}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{K+1}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}^{a}+e}\right],$
@@
$\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}^{a}+e}\in{\mathbb{R}}^{N}$, and $\mathbf{r}^{\prime}=[r_{1},r_{2},\ldots,r_{K+1}]\in{\mathbb{R}}^{K+1}$.
Since $(g_{j_{1}}({\bm{\Delta}}){\mathbf{X}})^{\top}(g_{j_{2}}({\bm{\Delta}}){\mathbf{X}})=0$ for all $j_{1},j_{2}\in[K+1]$, the matrix $\mathbf{H}^{\top}\mathbf{H}$ is a diagonal matrix with entries $\left\lVert\frac{g_{1}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{1}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}^{a}+e}\right\rVert_{2}^{2}$. Therefore, the singular values of $\mathbf{H}$ are
@@
$\sigma_{j}=\frac{\left\lVert g_{1}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}}{\left\lVert g_{1}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}^{a}+e}$
the right singular vectors are the standard basis elements $\mathbf{e}_{j}$ in ${\mathbb{R}}^{K+1}$, and the left singular vectors are
$\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}}.$
for $j\in[K+1]$. Hence, we have
$\left\lVert\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{\text{val, diag}}\right\rVert_{2}=\max_{j}\sigma_{j}$
@@
$\left\lVert\left(\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{\text{val, diag}}\right)^{-1}\right\rVert_{2}=\frac{1}{\min_{j}\sigma_{j}}.$
Suppose that $a=1$ and $e=0$. In this case, $\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{\text{val, diag}}$ is an isometry from the spectral domain to a subspace of the signal space ${\mathbb{R}}^{N}$. Analysis is the adjoint of synthesis. This analysis can be extended to the index parametrization case for diagonal synthesis, and to higher dimensional signals.
D.4 Uniform Approximation of Graphs with $J$ Eigenvectors
@@ D.4.1 Cut Norm
The cut norm has the following interpretation, which has precise formulation in terms of the weak regularity lemma [32, 65]. Any pair of (deterministic) graphs are close to each other in cut norm if and only if they can be described as pseudo-random graphs sampled from the same stochastic block model. Hence, the cut norm is a meaningful notion of graph similarity for practical graph machine learning, where graphs are noisy and can represent the same underlying phenomenon even if they have different sizes and topologies. In addition, the distance between non-isomorphic graphons is always positive in cut norm [66]. In this context, the work in [59] showed that GNNs with normalized sum aggregation cannot separate graphs that have zero distance in the cut norm. This means that the cut norm is sufficiently discriminative for practical machine learning on graphs.
D.4.2 The Constructive Weak Regularity Lemma in Hilbert Spaces
The following lemma, called the constructive weak regularity lemma in Hilbert spaces, was proven in [29]. It is an extension of the classical respective result from [65].
@@ D.4.2 The Constructive Weak Regularity Lemma in Hilbert Spaces
Lemma D.3.
[[29]] Let $\{\mathcal{K}_{j}\}_{j\in\mathbb{N}}$ be a sequence of nonempty subsets of a real Hilbert space $\mathcal{H}$ and let $\delta\geq 0$. Let $J>0$, let $R\geq 1$ such that $J/R\in\mathbb{N}$, and let $g\in\mathcal{H}$. Let $m$ be randomly uniformly sampled from $[J]$. Then, in probability $1-\frac{1}{R}$ (with respect to the choice of $m$), any vector of the form
@@
Let us use Lemma D.3, with $\mathcal{H}={\mathbb{R}}^{N\times N}$, and $\mathcal{K}_{j}=\mathcal{K}$ the set of symmetric rank one matrices of the form $\mathbf{v}\mathbf{v}^{\top}$ where $\mathbf{v}\in{\mathbb{R}}^{N}$ is a column vector. Denote by $\mathcal{Y}_{m}$ the space of linear combinations of $m$ elements of $\mathcal{K}$, which is the space of symmetric matrices of rank bounded by $m$. For the Hilbert space norm, we take the Frobenius norm. In the setting of the lemma, we take $g=\mathbf{M}$, and $g^{*}\in\mathcal{Y}_{m}$, and $\delta=0$. By the lemma, with probability $1-1/R$, any Frobenius minimizer $\mathbf{C}$, namely, that satisfies $\left\lVert\mathbf{M}-\mathbf{C}\right\rVert_{\rm F}=\min_{\mathbf{C}^{\prime}\in\mathcal{Y}_{m}}\left\lVert\mathbf{M}-\mathbf{C}^{\prime}\right\rVert_{\rm F}$, also satisfies
@@
$\leq$ $\left\lVert\frac{1}{2}\left\langle\bm{\mathbbm{1}}_{S\cup T}\bm{\mathbbm{1}}_{S\cup T}^{\top},\mathbf{M}-\mathbf{C}\right\rangle\right\rVert_{\rm F}+\left\lVert\frac{1}{2}\left\langle\bm{\mathbbm{1}}_{S}\bm{\mathbbm{1}}_{S}^{\top},\mathbf{M}-\mathbf{C}\right\rangle\right\rVert_{\rm F}+\left\lVert\frac{1}{2}\left\langle\bm{\mathbbm{1}}_{T}\bm{\mathbbm{1}}_{T}^{\top},\mathbf{M}-\mathbf{C}\right\rangle\right\rVert_{\rm F}$
@@
where for a set $S\subset[N]$, denote by $\bm{\mathbbm{1}}_{S}\in{\mathbb{R}}^{N}$ the column vector with $1$ at coordinates in $S$ and 00 otherwise.
@@
$\left\lVert\mathbf{M}-\mathbf{C}\right\rVert_{\square}\leq\frac{3\alpha}{2}\sqrt{\frac{R}{J}}.$
Lastly, note that by the best rank-$m$ approximation theorem (Eckart–Young–Mirsky Theorem [92, Thm. 5.9]), any Frobenius minimizer $\mathbf{C}$ is the projection upon the $m$ leading eigenvectors of $\mathbf{M}$ (or some choice of these eigenvectors in case of multiplicity higher than 1).
@@
Now, if we want to apply Thm. 4.1 to other GSOs ${\bm{\Delta}}$, we need to make some assumptions. Note that when the GSO is a Laplacian, we take as the leading eigenvectors the ones correspoding to the smallest eigenvalues, not the largest ones. To make the theorem applicable, we need to reorder the eigenvectors of ${\bm{\Delta}}$. This can be achieved by applying a decreasing function $h$ to ${\bm{\Delta}}$, such as $h({\bm{\Delta}})={\bm{\Delta}}^{-1}$. The role of $h$ is to amplify the eigenspaces of ${\bm{\Delta}}$ in which most of the energy of signals interest (the ones that often appear in the dataset) is concentrated. Under the assumption that $h({\bm{\Delta}})$ has entries bounded by some not-too-high $\alpha>0$, one can now justify approximating GSOs by low-rank approximations based on the smallest eigenvectors.
D.5 NLSFs on Random Geometric Graphs
@@
We follow a similar analysis to [74], mostly skipping the rigorous Monte-Carlo error rate proofs, as these are standard.
Let $\mathcal{S}$ be a compact metric space with metric ${\rm d}$ and with a Borel probability measure, that we formally call the standard measure $dx$. Let $r>0$, and denote by $B_{r}(x)$ the ball of radius $r$ about $x\in\mathcal{S}$. Let $V(x)$ be the volume of $B_{r}(x)$ with respect to the standard measure (note that $V(x)$ need not be constant). We consider an underlying Laplacian on the space $\mathcal{S}$ defined on signals $f:\mathcal{S}\rightarrow\mathbb{R}$ by
@@
where we assume $C=1$ in the analysis below without loss of generality. In case the integral in Eq. (22) is normalized by $1/r^{2}V(x)$, this Laplacian was called $\rho$-Laplacian in [15], which we denote by $\mathcal{L}_{r}$. Such a Laplacian is related to Korevaar-Schoen type energies [52], in which the limit case of the radius $r$ going to 0 is considered. It was shown in [15] that the $\rho$-Laplacian is self-adjoint with spectrum supported inside some interval $[0,Q]$, for some $Q>0$, where for some $0<R<Q$, the part of the spectrum in $[0,R)\cup(R,Q]$ is discrete (consisting of isolated eigenvalues with finite multiplicities). The intuition behind this result is that, for the $\rho$-Laplacian $\mathcal{L}_{r}$, we have
@@
The first term in the right-hand-side of Eq. (23) is a scaled version of the identity operator, and the second term is a compact self-adjoint integral operator, and hence has a discrete spectrum with accumulation point at 0, or no accumulation point. After showing that the sum of these two operators is self-adjoint, we end up with only one accumulation point of the spectrum of $\mathcal{L}_{r}$.
We show below that $\mathcal{L}$ is self-adjoint under the assumption that $V(x)$ is bounded from above. In this case it must also be positive semi-definite (the proof is equivalent to the positivity of the combinatorial graph Laplacian). Consider the decomposition
@@
The second term of Eq. (24) is a compact integral operator, and hence only has an accumulation point at 00, or has no accumulation point. The first term is a multiplication operator by the real-valued function $V(x)$, so it is self-adjoint and its spectrum is the closure of $\{V(x)\ |\ x\in\mathcal{S}\}$. We suppose that $V(x)$ is bounded from below by some $R^{\prime}>0$ and also bounded from above. This shows that $\mathcal{L}$ is bounded and self-adjoint (as a sum of two bounded self-adjoint operators). Under these assumptions, it is reasonable to further assume that the spectrum of $\mathcal{L}$ in the interval $[0,R)$, for some $R>0$, is discrete, with accumulation point at $R$. This happens, for example, if $V(x)=V$ is constant.
We now assume that the signal $f$ consists only of frequencies in $[0,R)$. This means that we can project $\mathcal{L}$ upon this part of the spectrum, giving a self-adjoint operator with discrete spectrum, and accumulation point of the eigenvalues only at $R$.
[…]
We consider the leading $J$ eigenvectors the ones corresponding to the smallest eigenvalues of $\mathcal{L}$ in the interval $[0,R)$, where $J\ll N$. We order the eigenvalues in increasing order. Let $\mathcal{PW}(J)$ be the space spanned by the $J$ leading eigenvectors of $\mathcal{L}$, called the Paley-Wiener space of $\mathcal{L}$. Let $p_{J}$ be the projection upon the Paley-Wiener space of $\mathcal{L}$. Let $R_{N}:L^{2}(\mathcal{S})\rightarrow L^{2}[N]$ be the sampling operator, defined by $(R_{N}f)_{j}=f(x_{j})$.
Let $f:\mathcal{S}\rightarrow{\mathbb{R}}$ be a bounded signal over the metric space and suppose that $f\in\mathcal{PW}(J)$. Denote $\mathbf{f}=(f(x_{n}))_{n=1}^{N}=R_{N}f$. By Monte Carlo theory, we have
[…]
So, in case the sampling is uniform. i.e., $w(x)=1$, we have
$\mathbf{L}\approx N\mathcal{L},$
pointwise. To make this precise, let us recall Hoeffding’s Inequality [43].
Theorem D.1 (Hoeffding’s Inequality [43]).
@@
$\mathbb{P}\left(\left\lvert\frac{1}{N}\sum_{n=1}^{N}(Y_{n}-\mathbb{E}[Y_{n}])\right\rvert\geq k\right)\leq 2\exp\left(-\frac{2k^{2}N}{(b-a)^{2}}\right).$
@@
$\left\lVert R_{N}\mathcal{L}f-\frac{1}{N}{\mathbf{L}}R_{N}f\right\rVert={\mathcal{O}}\left(J\sqrt{\frac{\log(1/p)+\log(N)+\log(2)}{N}}\right).$
@@
We now evoke the transferability theorem – a slight reformulation of Section 3 of Theorem 4 in [60].
Theorem D.2.
Let $\mathcal{L}$ be a compact operator on $L^{2}(\mathcal{S})$ and $\mathbf{L}$ an operator on $\mathbb{C}^{N}$. Let $J\in\mathbb{N}$ and let $R_{N}$ be the sampling operator defined above. Let $\mathcal{PW}(J)$ be the space spanned by the $J$ leading eigenvectors of $\mathcal{L}$, and let $f\in\mathcal{PW}(J)$. Let $g:{\mathbb{R}}\rightarrow{\mathbb{R}}$ be Lipschitz with Lipschitz constant $D$. Then,
$\left\lVert R_{N}g(\mathcal{L})f-\frac{1}{N}g({\mathbf{L}})R_{N}f\right\rVert\leq D\sqrt{J}\left\lVert R_{N}\mathcal{L}f-\frac{1}{N}{\mathbf{L}}R_{N}f\right\rVert.$
For every leading eigenvalue $\lambda_{j}$ of $\mathcal{L}$ let $\gamma_{i_{j}}$ be the leading eigenvalue of $\mathcal{L}$ closest to $\lambda$, where, if there are repeated eigenvalues, $i_{j}$ is chosen arbitrarily from the eigenvalues of $\mathbf{L}$ that best approximate $\lambda_{i}$. Let
$\delta=\min\{\alpha,\beta\},$
@@
$\alpha=\min_{j\in[J]}\min\big{\{}\ \min_{i_{j}\neq i\in[N]}\left\lvert\lambda_{j}-\gamma_{i}\right\rvert\ ,\ \min_{j\neq m\in[J]}\left\lvert\lambda_{j}-\lambda_{m}\right\rvert\ ,\ \min_{j\neq m\in[J]}\left\lvert\gamma_{i_{j}}-\lambda_{m}\right\rvert\ \big{\}}$
@@
$\beta=\min_{i\in[J],n\in[N]}\left\lvert\gamma_{i}-\gamma_{n}\right\rvert$
@@
$g_{j}(\mathcal{L})f=p_{j}^{\mathcal{L}}f,\quad g_{j}(\mathbf{L})R_{N}f=p_{i_{j}}^{\mathbf{L}}\mathbf{f},$
@@
Now, by the transferability theorem,
$\left\lVert R_{N}p^{\mathcal{L}}_{j}f-p^{\mathbf{L}}_{i_{j}}\mathbf{f}\right\rVert\leq D\sqrt{J}\left\lVert R_{N}\mathcal{L}f-\frac{1}{N}{\mathbf{L}}\mathbf{f}\right\rVert={\mathcal{O}}(\sqrt{J/N}).$
By induction over $j$, with base $j=1$, we must have $i_{j}=j$ for every $j\in[J]$.
@@
$\left\lvert\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}^{2}-\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}^{2}\right\rvert={\mathcal{O}}(J^{-1/2})$
@@
and $\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}+\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}$ is bounded from below by the constant $\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}$, we have
$\left\lvert\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}-\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}\right\rvert={\mathcal{O}}(J^{-1/2}).$
@@
Appendix E Additional Details on Experimental Study
We describe the experimental setups and additional details of our experiments in Sec. 5. The experiments are performed on NVIDIA DGX A100.
E.1 Semi-Supervised Node Classification